OpenStudy (anonymous):

proof that cos(inverse sinx) = squareroot(1-x^2 )

OpenStudy (anonymous):

Call theta = sin^-1 (x). sin (theta) = x. In a right triangle, the opposite side a = x, hypotenuse c = 1, thus b = sqrt (1 - x^2). Accordingly, cos (theta) = b/c = sqrt (1 - x^2)

OpenStudy (anonymous):

Thank you